# OpsEval: A Comprehensive IT Operations Benchmark Suite for Large Language Models

### v3

@@ 3.2. Quality Enhancement
The distribution of the questions across these eight tasks and three ability levels is depicted in Table 1.
Manual Review: In the manual review step, we asked Ops experts from the industry to inspect the results of the previous three automated steps, including confirming duplicate and invalid questions and examining the classification results of GPT-4. Experts were also asked to drop the questions unrelated to the Ops field. We split the dataset by n-folds and ensure each fold has at least two experts to review. By involving experts’ consensus, we ensure the reliability and authority of the dataset after automated processing.
Table 2. The number of questions in OpsEval, grouped by their scenarios.
Scenario Type Questions
Wired Network Multi-Choice 3901
5G Communication Multi-Choice 2615
Question-Answering 1162
Oracle Database Multi-Choice 497
Log Analysis Question-Answering 420
DevOps Question-Answering 154
Securities Information System Multi-Choice 91
Hybrid Cloud Multi-Choice 40
Financial IT Multi-Choice 40
Total Multi-Choice 7184
Question-Answering 1736
As listed in Table 2, this quality enhancement process resulted in a refined test set of approximately 7,000 multi-choice and 2000 question-answering questions.
3.3. Question Formatting
After quality enhancement of the collected questions, we perform question formatting for the upcoming evaluation. Each question is structured to include the context, answer, task, and ability. Figure 2 shows three examples of the formatted questions. This structured approach facilitates easy evaluation and analysis.
Figure 2. Three examples of the processed questions
3.4. Evaluation Settings
In our evaluation of LLMs within the Ops domain, we categorize our assessment questions as multi-choice and question-answering.
Multi-choice questions offer a structured approach with definitive answers. These questions are straightforward and provide a clear metric for assessment. However, given the intricacies of advanced models, they may be influenced by the options provided, leading to their responses being driven more by pattern recognition rather than a proper understanding of the content.
We use accuracy as the metric. The output of LLM may contain additional strings besides the chosen option, even if we make explicit requests and use few-shot in the prompt. Therefore, we use a choice-extracting function based on regular expressions to extract the predicted answer of LLMs. Then, we calculate the accuracy based on the extracted answer and the ground-truth labels.
Question-answering questions do not come with predefined options. It necessitates the model to rely more on its comprehension and knowledge base, offering a clearer insight into its cognitive capabilities. Such questions can better assess LLMs’ ability to generate coherent and contextually relevant responses.
We use two metrics for question-answering questions: one is based on word overlaps, and the other is based on semantic similarity. For the first type, we use ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002), widely used in NLP tasks, especially in the translation task. For the second type, we use GPT-4 and experts to obtain the output score of LLMs, called GPT4-Score and Expert-Evaluation, designed explicitly in OpsEval. We provide a more detailed description of these metrics in Appendix A.5.
For GPT4-Score, we provide GPT-4 with the question and the reference answer, followed by one anonymous model’s output. Figure 2(b) shows the prompt we used to ask GPT-4 to score.
We design expert evaluation for manually scoring LLMs’ outputs based on three criteria highly related to Ops’ needs. The three criteria in consideration are as follows:
@@ 3.4. Evaluation Settings
For each output to a question of an LLM, we asked experts to score it between 0 and 3 for each criterion. During the scoring, the raw question, the detailed answer and its key points, and the output of an anonymous model are given at each iteration. Since avoiding bias towards specific models is important, no information about the model is given.
Fairness Consideration We are currently making 20% of the data available to the public for Ops community contribution and research purposes, yet for fairness of the evaluation, the complete version of OpsEval dataset is kept private and not disclosed. To evaluate a new model, users can submit a Docker image with an initialization script when starting a container based on it. We will run the evaluation automatically and obtain the result on the OpsEval website. Users can choose to disclose their results on the leaderboard of OpsEval or not based on their preference.
@@
Figure 3. Prompts for ChatGPT used in the framework of OpsEval.
4. Experiment Design
In this section, we will show the experiment design of OpsEval. We evaluate various LLMs on OpsEval, aiming to understand the multiple abilities of different LLMs in addressing different question types (multi-choice and question-answering questions) and tasks. We also evaluate LLMs with different quantization parameters.
4.1. Models
Table 3. Models evaluated in this paper
Model Creator #Parameters Access [^3] The “access” column in the table shows whether we have full access to the model weights or can only access them through API.
GPT-4 OpenAI undisclosed API
GPT-3.5-turbo OpenAI undisclosed API
ERNIE-Bot-4.0 Baidu undisclosed API
GLM4 Tsinghua Zhipu undisclosed API
GLM3-turbo Tsinghua Zhipu undisclosed API
LLaMA-2 Meta 7/13/70B Weights
Qwen-Chat Alibaba Cloud 7/14/72B Weights
InternLM2-Chat Shanghai AI Laboratory 7/20B Weights
DevOps-Model-Chat CodeFuse 14B Weights
Baichuan2-Chat Baichuan Intelligence 13B Weights
ChatGLM3 Tsinghua Zhipu 6B Weights
Mistral Mistral 7B Weights
Table 4. GPTQ models for LLaMA-2-70B
Model Size #GPTQ Dataset Disc
LLaMA-2-70B 140GB /
Raw LLaMA-2-70B model.
LLaMA-2-70B-Int4 35.33GB wikitext
4-bit quantization model.
LLaMA-2-70B-Int3 26.78GB wikitext
3-bit quantization model.
As shown in Table 3, we evaluate popular LLMs covering different weights from different organizations. The detailed information of all LLMs in Table 3 can be found in Appendix B.1.
Besides, we evaluate LLaMA-2-70B with multiple quantization parameters to get an overview of the effect of different quantization parameters. Specifically, we use GPTQ (Frantar et al., 2022) models with 3-bit and 4-bit quantization parameters. [^4] The two quantization models are downloaded from https://huggingface.co/TheBloke/Llama-2-70B-chat-GPTQ. The size of LLaMA-2-70B is calculated based on 70B parameters. The two GPTQ models evaluated in our experiments are calibrated on wikitext (Merity et al., 2016), a language modeling dataset extracted from the set of verified Good and Featured articles on Wikipedia. GPTQ is a post-training quantization (PTQ) method to make the model smaller with a calibration dataset. It is a one-shot weight quantization method based on approximate, highly accurate, and efficient second-order information. The details of GPTQ models can be found in Table 4.
4.2. Prompting Techniques
@@
To get a comprehensive overview of the performance of popular LLMs on OpsEval, we use various settings to perform the evaluation. We evaluate LLMs in zero and few-shot settings (3-shot in our implementation). For each setting, we evaluate LLMs in four sub-settings of prompt engineering, that is, naive answers (Naive), self-consistency (SC (Wang et al., 2023b)), chain-of-thought (CoT (Wei et al., 2023a)), self-consistency with chain-of-thought (CoT+SC). We design prompts for both English and Chinese languages.
Naive. The Naive setting is to expect the LLMs to generate the answer without any other explanations. Since we have the task type of each question, we integrate the task into the prompt.
SC. Self-consistency is selecting the most consistent answer among several queries on LLMs. Although it aims “to replace the naive greedy decoding used in chain-of-thought prompting,” it can generate naive answers as it may generate different answers with the same prompts. We set the number of queries in SC to 5.
CoT. The CoT setting aims to enable LLMs to obtain complex reasoning capabilities through intermediate reasoning steps. We construct specific prompts for CoT setting in both zero and few-shot evaluations. Details of the prompt construction can be found in Appendix B.2.
CoT+SC. We combine CoT and SC to boost the performance of CoT prompting. By SC, We choose the consistent reasoning path and answer several of the same queries. Like the SC setting, we set the number of queries in CoT+SC to 5.
As for question-answering problems, we combine each question’s task type and ability and the question as the prompt for LLMs. Figure 9 in Appendix A.4 shows an example of constructing the prompt.
5. Evaluation
5.1. Benchmark Leakage Test
As a benchmark suited for LLM, it is necessary to ensure that the test set has not been included in the model’s pretraining process. Otherwise, the potential bias may damage the fairness of the benchmark. We adapted the methodology from Skywork (Wei et al., 2023b) to perform a leakage test on OpsEval’s dataset. We evaluate the LLM loss on samples (a sample is a concatenation of question and answer) from different datasets for several foundation models and calculate the average loss. For each dataset, we compare LM loss on the test split ($L_{test}$) and a specially curated reference set ($L_{ref}$ ) generated by GPT-4, designed to mimic the testing dataset. We define a key metric: $\Delta L=L_{test}-L_{ref}$, serving as an indicator of potential test data leakage during the training of the LLM, i.e., a lower value suggests possible leakage. Notice that while Skywork (Wei et al., 2023b) only asked GPT-4 to generate similar questions to GSM8K, we require GPT-4 to rewrite the question while preserving its original meaning and accuracy. By preserving the original meaning, if the $\Delta L$ of the test set is a lower value, it suggests that the LLM’s lower $L_{test}$ originates from over-fitting the distribution of tokens in the test set rather than internalizing the knowledge behind the questions, thereby suggesting a leakage in the test set.
Table 5. Measurement of potential test data leakage during the training of LLM. This demonstrates the unbiased nature and non-leakage of the OpsEval test set.
Dataset $L_{test}$ $L_{ref}$ $\Delta L$
Alpaca 1.994033 2.354260 -0.360228
Alpaca-GPT4 1.498862 1.763663 -0.391062
CEval 2.570809 2.309943 0.260866
MMLU 2.547598 2.189870 0.357728
OpsEval 1.885437 1.728079 0.105095
Table 5 shows the results of leakage measurement. In addition to the two standard evaluation benchmarks (CEval (Huang et al., 2023) and MMLU (Hendrycks et al., 2021)), we conducted the same experiments on the alpaca dataset (Taori et al., 2023) and the Alpaca-GPT4 dataset (Peng et al., 2023), which is likely used in the pre-training of large models, using its $\Delta L$ as reference. The corpora likely involved in training show a significantly smaller $\Delta L$, whereas the loss for the OpsEval dataset remains at a relatively small positive value. This demonstrates the unbiased nature and non-leakage of the OpsEval test set. The models we used in the leakage test are listed in Appendix B.1.
5.2. Overall Performance
@@
Table 6. LLMs’ overall performance on Wired Network Operations English test set (3-shot). Models are ranked based on their best performance (marked as bold) among different settings.
Model Naive SC CoT CoT+SC
GPT-4 / / 88.70 /
GLM-4 64.77 64.77 77.06 77.06
GPT-3.5-turbo 68.30 68.30 70.90 72.50
Qwen-72B-Chat 70.32 70.32 70.13 70.22
ERNIE-Bot-4.0 60.00 60.00 70.00 70.00
LLaMA-2-70B 55.00 56.20 66.80 67.20
DevOps-Model-14B-Chat 63.85 61.96 41.15 44.01
GLM-3-turbo 59.53 59.53 63.65 63.65
Qwen-14B-Chat 62.60 59.70 50.58 55.88
LLaMA-2-13B 53.30 53.00 56.80 61.00
InternLM2-Chat-20B 60.48 60.48 45.10 45.10
LLaMA-2-7B 48.20 46.80 52.00 55.20
Qwen-7B-Chat 52.10 51.00 48.30 49.80
Baichuan2-13B-Chat 51.90 51.60 44.50 47.45
InternLM2-Chat-7B 48.2 48.2 49.74 49.74
Mistral-7B 47.22 47.22 45.58 45.58
ChatGLM3-6B 42.10 42.10 43.47 43.47
The results of the few-shot evaluation with four settings on the Wired Network Operation test set are shown in Table 6. Results of the other scenarios and settings are shown in Appendix B.3.  [^5] Due to the consideration of time, cost, and API rate limits, for GPT-4, we only make the 3-shot evaluation with the CoT setting to serve as an upper bound of all LLMs to provide a reference. From the overall performance results, we can come to several findings.
On both English and Chinese questions, GPT-4 consistently outperforms all other models, surpassing the best performances of all other LLMs. LLM with larger parameter size like Qwen-72B-Chat, ERNIE-Bot-4.0, and LLaMA-2-70B, when employing the Self-Consistency and Chain of Thought (CoT) prompt methods, approach the performance of ChatGPT. Smaller models, such as DevOps-Model-14B-Chat and Qwen-14B-Chat, exhibit competitive performance in multi-choice questions, approaching the capabilities of models with 70B parameters, thanks to their fine-tuning process and the quality of their training data.
Furthermore, the effectiveness of the four prompt settings varies across different LLMs. We examine LLMs’ zero-shot and few-shot performances under four settings mentioned earlier in Sec. 4 for both English and Chinese test sets. From the evaluation results, we can conclude the following observations:
  1. (1)
    For most models, the performance improves from the setting of Naive to SC, CoT, and SC+CoT. Notably, few-shot performance is better than zero-shot performance.
  2. (2)
    Among these settings, CoT prompts yield the most significant improvement in LLMs’ answering capability. SC prompts result in relatively minor improvements. LLMs’ responses are consistent across repeated questions, aligning with the desired outcome in operational tasks where reliability and consistency are essential.
  3. (3)
    LLMs fine-tuned specifically for Chinese exhibit better performance on English and Chinese test sets than LLMs without Chinese fine-tuning. We discuss further insights into these observations in Appendix. B.5.
  4. (4)
    In a few cases, more advanced evaluation methods surprisingly lead to poorer results. The detailed analysis for this can be found in Appendix B.6.1.
5.3. Performance on Different Tasks and Abilities
To investigate how LLMs perform in each Ops task and to what extent they possess the abilities of Knowledge Recall, Analytical Thinking, and Practical Application, we summarize the result of different parameter-size We group LLMs’ performance based on the task and ability classification mentioned in Sec. 3.2 and plot them on two radar charts in Figure 4.
Figure 4. LLMs’ performance on eight tasks and three abilities
In terms of the eight tasks we tested, LLMs generally yield higher accuracy in General Knowledge tasks, while their performance drops and varies drastically in highly specialized tasks like Automation Scripts and Network Configuration, reflecting the impact of specialized corpus and domain knowledge on the performance of LLMs.
Among the three abilities, LLMs perform the best in Practical Application, followed by Knowledge Recall. It is foreseen that LLMs perform poorly in Analytical Thinking questions, as accurately deducing conclusions from existing facts for LLMs remains a challenging research topic. LLMs perform best in Practical Application because the LLMs we test are trained on the corpus where best Ops practices are involved, familiarizing the LLMs with solutions to many real-world tasks.
By grouping LLMs by their parameter size, we find that although LLMs with 10B-20B parameters have higher accuracy in their best cases compared with LLMs with no more than 7B parameters, different 10B-20B LLMs’ performance varies drastically, sometimes even lower than that of 7B. LLMs with no more than 7B parameters, on the other hand, have a more stable performance range within the group.
5.4. Performance on Question-Answering
@@
Table 7. LLMs’ performance on English network operations question-answering problems. GPT4 for GPT4-Score, FL for Fluency, AC for Accuracy, EV for Evidence, Total is the sum of the previous three columns.
Model ROUGE BLEU GPT4 Expert Evaluation
FL AC EV Total
@@
Figure 5. Scatter plot and trendline of GPT4-Score compared to Expert Evaluation score.
Table 8. Correlation coefficients between Expert-Evaluation Total and other metrics
Metric GPT4-Score BLEU-Score ROUGELsum
Correlation coefficient 0.9211 0.6108 -0.4559
Table 9. Correlation coefficients between GPT4-Score and sub-metrics of Expert-Evaluation
Metric Fluency Accuracy Evidence Total
Correlation coefficient 0.8700 0.9084 0.7978 0.9211
Table 7 presents the evaluation results of 200 question-answering English questions across four metrics: ROUGE, BLEU, GPT4-Score, and Expert-Evaluation, sorted by GPT4-Score results. We have conducted question-answering tests on GPT-4 to verify its capabilities before using it for scoring. Since GPT-4 is currently the most capable among large models, its performance on question-answering questions surpasses other models. However, we did not include it in the table to avoid bias and misunderstanding.
The rankings based on ROUGE and BLEU scores do not align well with GPT4-Score and Expert-Evaluation, as shown in table 8. LLMs with poor performance may generate keywords, resulting in higher ROUGE and BLEU scores. In contrast, LLMs with good performance might receive lower ROUGE/BLEU scores due to differences in wording compared to the standard answers.
Regarding GPT4-Score, the rankings closely resemble those based on Expert-Evaluation. In table 9, we calculate the correlation coefficients between GPT4-Score and different sub-metrics of Expert-Evaluation to gain more insights. Among the three metrics, rankings of GPT4-Score align most closely with the Accuracy metric, suggesting that GPT4 is most reliable on the factuality with its vast knowledge base. The format and length of the generated content also heavily influence GPT4-Score, as suggested by the high positive correlation between GPT4-Score and Fluency. On the other hand, there are more discrepancies in rankings concerning the Evidence metric, indicating that GPT4-Score needs to consider the role of arguments and evidence in cases where answers are ambiguous.
In the Expert-Evaluation, where Evidence is a significant criterion, LLMs with more elaborate arguments can outperform GPT-3.5-turbo in total scores even when their Accuracy scores are much lower than the latter.
5.5. Performance on Different Quantization parameters
Figure 6. LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted using the naive setting on the English Wired Network Operations test set.
Figure 6 shows the accuracy of LLaMA-2-70B of different quantization parameters on English multi-choice questions. We do both few-shot and zero-shot evaluations with the naive setting. Using quantization during inference brings a performance degradation.
LLaMA2-70B-Int4 can achieve an accuracy close to LLaMA-2-70B without quantization. Specifically, on English multi-choice questions, the accuracy of the GPTQ model with 4-bit quantization parameters is 3.50% lower in zero-shot evaluation and 0.27% in few-shot evaluation compared to LLaMA-2-70B. As for Chinese questions, the accuracy of LLaMA2-70B-Int4 is 3.67% lower in zero-shot evaluation and 5.18% in few-shot evaluation compared to LLaMA-2-70B.
However, LLaMA2-70B-Int3 has a performance degradation that cannot be ignored, as shown in Figure 6. On average, the accuracy of LLaMA2-70B-Int3 has a 12.46% degradation compared to LLaMA-2-70B and a 9.30% degradation compared to LLaMA2-70B-Int4. The reason may be that the information of the full-sized model is lost too much in 3-bit quantization.
6. Discussion
6.1. Implications of the evaluation results
We summarize our key observations from the evaluation here:
  • Few-shot and CoT can significantly increase performance if the model is tuned to adapt to these techniques, while SC may have little influence on highly consistent LLMs.
  • LLMs perform best in general knowledge while deducing specialized conclusions remains challenging.
@@ 6.1. Implications of the evaluation results
  • GPT4-Score is suitable as an automatic metric in large-scale qualitative evaluations.

### v4

@@ 3.2 Preprocessing
Manual Review: In the manual review step, we asked Ops experts from the industry to inspect the results of the previous three automated steps, including confirming duplicate and invalid questions and examining the classification results of GPT-4. In our work, an expert is defined as an individual with ten or more years of professional experience in their field, whether as an employee or a researcher. Experts were also asked to drop the questions unrelated to the Ops field. We split the dataset by n-folds and ensure each fold has at least two experts to review. As listed in Table 2, this quality enhancement process resulted in a refined test set of approximately 7,000 multi-choice and 2,000 question-answering questions.
3.3 Evaluation Settings
Multi-choice questions offer a structured approach with definitive answers. These questions are straightforward and provide a clear metric for assessment. We use accuracy as the metric. A choice-extracting function based on regular expressions is used to extract the predicted answer of LLMs. Then, we calculate the accuracy based on the extracted answer and the ground-truth labels.
Question-answering questions do not come with predefined options. We use two metrics for question-answering questions: one is based on word overlaps, and the other is based on semantic similarity. For the first type, we use ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002). For the second type, we use LLM judges and human experts to evaluate the outputs of LLMs, called FAE-Score and Expert-Evaluation, designed explicitly in OpsEval. For these two metrics, we design three criteria highly related to Ops’ needs. The three criteria in consideration are as follows:
@@ 3.3 Evaluation Settings
For FAE-Score, we provide the judge model with the question and the reference answer, followed by one anonymous model’s output. We use GPT-4 as the judge model of FAE-Score based on evidence that GPT-4 can reliably match human preferences with over 80% agreement (Zheng et al., 2023). In the subsequent validation section §5, we also verified the consistency between GPT-4’s scores and expert scores. Figure 2(b) shows the prompt we used to ask the judge model to score. For Expert-Evaluation, we asked experts to score it between 0 and 3 for each criterion. During the scoring, the raw question, the detailed answer and its key points, and the output of an anonymous model are given at each iteration. For the demographical information of the experts, see Appendix A.1.
@@
Figure 2: Prompts for ChatGPT used in the framework of OpsEval.
@@
Prompting Techniques. We use various settings to evaluate LLMs on OpsEval to get a comprehensive overview of their performance. We evaluate LLMs in zero and few-shot (3-shot) settings. For each setting, we evaluate LLMs in four sub-settings of prompt engineering, that is, naive answers (Naive), self-consistency (SC) (Wang et al., 2023b), chain-of-thought (CoT) (Wei et al., 2023), self-consistency with chain-of-thought (CoT+SC). We set the number of queries in SC to 5.
Models. We evaluate popular LLMs covering different weights from different organizations. The model selection was guided by specific criteria: We aimed to include the latest and most advanced large language models, with a particular focus on those capable of handling Chinese input. The detailed information of all 21 LLMs can be found in Table 9 in Appendix B.1.
Fairness Consideration We are currently making 20% of the data available to the public for Ops community contribution and research purposes, yet for fairness of the evaluation, the complete version of the OpsEval dataset is kept undisclosed. To evaluate a new model, users can submit a Docker image with an initialization script when starting a container based on it. We will run the evaluation automatically and obtain the result on the OpsEval website. Users can choose to disclose their results on the leaderboard of OpsEval or not based on their preference.
This released 20% dataset serves as sample questions so that the researchers will know the types and topics of questions and answers expected by the benchmarks. This allow researchers to gain intuition and insights on how to improve their models. Model developers can use the 20% dataset for locally evaluating by themselves their model’s performance, enabling faster iterations of their model training. The 20% dataset can be used as a seed to generate QA pairs using automatic QA generation algorithms (Wang et al., 2023c), ultimately providing more Ops text data for improving new models.
API Exposure. For large language models, user requests constitute private data that should be carefully protected. We believe that ensuring data privacy and not using such data for subsequent training is a fundamental requirement that all publicly accessible models must meet. The APIs we utilized for model evaluation explicitly guarantee that user data will not be used for model training (Microsoft, 2023; BigModel, 2023; Cloud, 2023). When evaluating closed-source models such as ChatGPT and GLM, the questions in this non-disclosed questions were sent (thus exposed) to their APIs, but the answers to these questions were still non-disclosed.
4 Result Analysis
4.1 Overall Performance
@@
Figure 3: LLMs’ overall performance on Wired Network Operations English test set (3-shot). Models are ranked based on their mean accuracy among different settings. The error bars represent the variance in the model’s accuracy across different prompting techniques.
Table 4: LLMs’ overall performance (Accuracy%) on Wired Network Operations English test set (3-shot). Models are ranked based on their best performance among different settings.
Model Naive SC CoT CoT+SC Best Score Variance
GPT-4 / / 88.70 / 88.70 /
Claude-3-Opus 79.70 78.36 81.24 82.28 82.28 2.9638
Qwen2-72B-Instruct 76.91 76.91 80.12 80.65 80.65 4.0720
Qwen2-7B-Instruct 65.74 65.74 67.42 67.42 67.42 0.9408
GLM-4 64.77 64.77 77.06 77.06 77.06 50.3480
GPT-3.5-turbo 68.30 68.30 70.90 72.50 72.50 4.2800
Qwen-72B-Chat 70.32 70.32 70.13 70.22 70.32 0.0083
ERNIE-Bot-4.0 60.00 60.00 70.00 70.00 70.00 33.3333
LLaMA-2-70B 55.00 56.20 66.80 67.20 67.20 43.5866
Qwen1.5-14B-Chat 52.23 53.52 59.53 64.17 64.17 23.0368
DevOps-Model-14B-Chat 63.85 61.96 41.15 44.01 63.85 139.6604
GLM-3-turbo 59.53 59.53 63.65 63.65 63.65 5.6581
Qwen-14B-Chat 62.60 59.70 50.58 55.88 62.60 26.9921
Meta-LLaMA-3-8B 41.03 42.07 62.45 62.62 62.62 110.2314
LLaMA-2-13B 53.30 53.00 56.80 61.00 61.00 13.9758
InternLM2-Chat-20B 60.48 60.48 45.10 45.10 60.48 78.8481
LLaMA-2-7B 48.20 46.80 52.00 55.20 55.20 14.4366
Qwen-7B-Chat 52.10 51.00 48.30 49.80 52.10 2.6600
Baichuan2-13B-Chat 51.90 51.60 44.50 47.45 51.90 12.5822
Gemma-7B 30.24 30.24 51.56 51.56 51.56 151.5141
InternLM2-Chat-7B 48.2 48.2 49.74 49.74 49.74 0.7905
Mistral-7B 47.22 47.22 45.58 45.58 47.22 0.8965
ChatGLM3-6B 42.10 42.10 43.47 43.47 43.47 0.6256
Gemma-2B 26.63 26.63 37.54 37.54 37.54 39.6760
The results of the few-shot evaluation with four settings on the Wired Network Operation test set are shown in Figure 3. Results of the other sub-domains and settings are shown in Appendix B.4.  [^6] Due to the consideration of time, cost, and API rate limits, for GPT-4, we only make the 3-shot evaluation with the CoT setting to serve as an upper bound of all LLMs to provide a reference. Overall, open-sourced LLMs yield evaluation results on the OpsEval benchmark generally worse than those in general domains like MMLU (Hendrycks et al., 2021) and CEval (Huang et al., 2023). This comparison highlights the necessity of explicitly fine-tuning OpsLLM for the Ops field. Closed source models like GPT-4 and GLM-4 consistently outperform open source models, while smaller models, such as Qwen1.5-14B-Chat, exhibit competitive performance in multi-choice questions, thanks to their fine-tuning process and the quality of their training data. However, their large variance across the four settings suggests that such models may have worse robustness under different prompts. Furthermore, we observed significant variability in how different LLMs respond to various prompt engineering techniques. Given the critical importance of stability in the Ops domain, it is essential to consider a model’s sensitivity to prompts when selecting foundation model. Further research into prompt engineering is needed to improve model performance and reliability in this domain.
Observations: 1) Few-shot and CoT can significantly increase performance if the model is tuned to adapt to these techniques, while SC may have little influence on highly consistent LLMs. 2) Smaller models with weaker natural language abilities are less stable with advanced prompts. Simpler prompts work better for them.
Pratical Lesson: The choice of fundamental models should be a balance between their performance (average score) and robustness (variance) under different prompt settings.
4.2 Performance on Different Tasks and Abilities
Figure 4: LLMs’ performance on eight Ops sub-domains, eight tasks and three abilities. Each colored area presents the lower and upper bound of the corresponding parameter-size group.
Figure 5: Heatmap of failure case distribution regarding tasks and abilities. The values represent the proportion of failure cases across all LLMs; redder areas indicate higher failure rates.
To investigate how LLMs perform in each Ops sub-domain and each task, and to what extent they possess the general abilities, we summarize the result of different parameter-size groups of LLM and plot them on three radar charts in Figure 4. Regarding the eight tasks we tested, LLMs yield higher accuracy in General Knowledge tasks, while their performance drops and varies drastically in highly specialized tasks like Automation Scripts and Network Configuration, reflecting the impact of specialized corpus and domain knowledge on the performance of LLMs. By grouping LLMs by their parameter size, we find that although LLMs with 10B-30B parameters have higher accuracy in their best cases compared with LLMs with no more than 10B parameters, different 10B-20B LLMs’ performance varies drastically. To provide systematic practical lessons for researchers in the operations domain on pre-training and fine-tuning OpsLLM, we have analyzed the error rates of LLMs across the 8 tasks and 3 abilities in Figure 5. By examining the focus areas across different categories, we have identified key research targets for capability training.
Observations: Among the 24 categories of results, models performed the worst in Analytical Thinking for Automation Scripts. This indicates that current models can only recall the learned scripts but struggle to infer their logical relationships. Similarly, Analytical Thinking showed the lowest performance across the three major tasks, indicating that current OpsLLM models still have some way to go before becoming foundational models for Ops Agents. Thus, researchers should focus on inference-related SFT (supervised fine-tuning) datasets.
Insights: 1) Among different sub-domains of Ops, 5G communication and database demand further pretraining and fine-tuning of LLM. 2) To be capable of an Ops agent, the foundation model must be able to make a connection between specialized domain knowledge.
4.3 Performance on Question-Answering
@@
Table 5: LLMs’ performance on English network operations question-answering problems. Total is the sum of the previous three columns.
Model ROUGE BLEU FAE-Score Expert Evaluation
Fluency Accuracy Evidence Total
@@
Table 5 presents the evaluation results of 200 question-answering English questions across four metrics: ROUGE, BLEU, FAE-Score, and Expert-Evaluation, sorted by FAE-Score results. To gain more insight into how different metrics perform in QA evaluation, we use Figure 16 (see in Appendix B.8.2) as a case analysis. While BLEU and ROUGE are efficient in natural language comparison, they lack semantic information to determine which part of the context is more important than others. Knowing that a given benchmark evaluates QA based on BLEU/ROUGE, there is an obvious way to trick the metric: repeat patterns occurring in the question, gaining a higher possibility to match some patterns in the reference answer. Due to their lack of semantic information related to Ops and the potential hack, traditional metrics like BLEU are unsuitable for specialized benchmarks. Instead, with specialized prompting, FAE-Score can pick up those important keywords and not be influenced by repeated words that contain no useful information. In a later section, we discuss the alignment between different metrics and expert evaluation, validating the effectiveness of FAE-Score in automated QA evaluation within the Ops domain.
@@
Practical Lesson: FAE-Score is suitable for large-scale qualitative evaluations in the Ops field.
4.4 Performance on Different Quantization parameters
We conducted experiments on different quantized versions of LLaMA-2-70B and obtained various results and conclusions [^7] For detailed results, please see Appendix B.5.. Overall, although the performance of the INT4 version decreases in both English and Chinese, the decline does not exceed 10%. However, the performance drop in the INT3 version is more significant, requiring careful consideration in practical applications.